what is the net work done on a body moving in a non uniform circular motion with
atagential = bt(b is positive constant

Grade:12

1 Answers

Vikas TU
7 years ago
at = dw/dt = bt^2 => dv = bt^2 dt
integerating both sides,
w = bt^3/3
and also,
dthetha/dt = bt^3/3
thetha = bt^4/12
Work = F x displace,ent that is:
 = m*at*thetha
out the values.
where m is the mass of the body.